OpenStudy (anonymous):

Could I find the two missing angle measures if I know some of the side lengths of a right triangle? How would I do this?

OpenStudy (anonymous):

yes... you use the trigonometric function such as sin(x) = opp/hyp cos(x) = adj/ hyp, or tan(x) = opp/adj when 1 side and 1 angle is given.. You use the Pythagorean theorem when 2 sides of the right triangle is given..

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (anonymous):

You're welcome !